Sol Summary – January 4th

Crew 186 Sol 4 Summary Report 04JAN2018

Sol 4

Summary Title: The Moons of Mars and the Greens of Earth

Author’s name: Max Fagin

Mission Status: All nominal

Sol Activity Summary: Geologically, Earth’s moon has very little in common with Mars. The geological processes that have shaped the surface of our Moon (vulcanism and impacts) are very different than the geological processes that have shaped the surface of Mars (water and wind). But sitting in an intriguing middle ground are Mars’ moons, Phobos and Deimos. By most metrics, Mars’ moons are fairly typical asteroids, but their proximity to Mars has caused their surfaces to become blanketed in Martian material kicked up by asteroid strikes and swept up by the moons as they snowplow through the space surrounding the planet. As three members of our crew have done thesis level work on how hard it is to land on Mars, we are acutely aware that the first Mars sample return mission may not sample Mars at all, but rather, the impact debris that Mars has deposited onto its moons. In this theme, today a team composed of Cesare, Melanie, Sam and Mark conducted an EVA to the region 5 km North West of the Hab known as “The Moons” (Yellow Moon, Beige Moon, Gray Moon and White Moon). Last night, Cesare gave the entire crew a seminar on the geology of the surrounding area so we could be on the lookout for interesting features on geology EVAs like this one.

We had originally budgeted time for three stops in the region for spectral measurements and sample collection, but the process of strapping the expensive and delicate equipment to our ATVs meant our estimates were incorrect, and we reached our turnaround time after only sampling one. So we must add this site to the increasingly long list of sites that will get a follow-up in a later EVA. Far more science to be done than scientists to do it will be part of a real mission to Mars, but it doesn’t make that constraint any easier to accept!

When the EVA team returned, they were greeted by the sight of the NorCal Mars Society rover wheeling out to meet their ATVs. The rover was successfully navigated for three complete circuits of the campus while the team was out. The ideal Mars exploration mission will utilize robotic resources of every kind (orbital, airborne, and surface) and it’s good to have at least one of those joining our mission.

This evening, Mark took us on a tour of GreenHab where the microgreens he planted 2 days ago are already starting to sprout. Flowers from GreenHab already adorn our dinner table, and near the end of our mission, we expect GreenHab’s food will too.

Look Ahead Plan: In an attempt to check off our science goals more rapidly, tomorrow’s EVA will be a hybrid geology and navigation EVA along the southern end of Cow Dung Rd. The geology team will explore sites south of kissing camel ridge, and drop off the navigation team to attempt a walk back under limited visibility conditions with the navigation beacon.

Anomalies in work: ATV 300 wouldn’t start this morning. We tried starting it again after returning from EVA, after it had been sitting in the sun for 2 hours, and it started up as expected. It was probably just too cold.

Weather: 0C, Winds Calm, Visibility 10 miles, Barometer 30.30 inHg

Crew Physical Status: Healthy

EVA: The Moons, 2.5 hours, geological sampling and spectra of clays in the region

Reports to be filed: Sol Summary, Operations Report, GreenHab Report, Science Report (Geology), Journalist Report w. Photo of the Day, EVA Report,EVA Request

Support Requested: We are exceeding our maximum daily water consumption target of 40 gal/sol as GreenHab operations ramp up, and are implementing tighter controls on water use to allow GreenHab as much water as possible. We are considering implementing a “3 bin” dishwashing system, any other recommendations on effective water saving strategies would be welcome.

Operations Report – January 4th

Crew 186 Operations Report 04JAN2018

SOL: 04

Name of person filing report: M. Grande

Non-nominal systems: None

Notes on non-nominal systems: Generator system limping along again on the now-nominal routine. It was, however, displaying “Low Battery” and running at only 5.6 V this morning, running on the old 10/17 battery which had been fully charged. I will now be running the 10/17 battery every evening and charging that one during the day, just as we were doing before while keeping the new 9/17 battery in storage (fully charged).

Generator (hours run): 14.1h

Generator turned off, charging battery at 8:40am

Generator turned on at 5:10pm

Solar— SOC

@ 8:40am : 100%

@ 5:10pm : 88%

Diesel: 58%

Propane: 35%

Ethanol Free Gasoline (5 Gallon containers for ATV): 8.75 Gallons

Water (trailer): 10 Gallons

Water (static): 351 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 129033.7 Gallons

Toilet tank emptied: No

ATVs Used: Honda, 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: 0.25 Gallons

# Hours the ATVs were used today: 01:00 hours

Notes on ATVs: ATVs were nominal except for 300, whose engine I couldn’t keep running for the life of me. I tried various combinations and levels and durations of choking and throttling, with input from the crew. The engine just kept dying on me immediately as soon as I let off the gas. Two hours later, GreenHab Officer Mark was able to turn it on immediately to bring it slightly back to its parking space.

Deimos rover used: No

Hours: 97.4

Beginning charge: 100%

Ending charge:

Currently charging: No

Sojourner rover used: Assigned to director only.

Hours: 5.1

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Spirit rover used: No

Hours: 8.7

Beginning charge: 100%

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: 4.3

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours: 3.1

Beginning charge: 100%

Ending charge:

Currently charging: No

HabCar used and why, where? No

General notes and comments: So much fun riding the ATVs out all together today!! We took an ATV-only road out to the Martian moons, over twisty and bumpy roads. We hiked only a little bit around Yellow Moon and Beige Moon, and I manned the GoPro for some video footage. We’re all pretty comfortable on the ATVs now, so it was a nice ride. Today’s challenge, though, was implementing more bungee cords and also Commander Max’s webbing/straps to secure down the big Trek (portable spectrometer) case plus a tool box and emergency water. Rather difficult in bulky EVA gloves!

Other difficult things: seeing inside the gas tank with that bubble helmet. Basically impossible, so I had to search for the glisten of the gas inside. Also, those large darned rubber gloves. They’re even worse than the bulky EVA gloves, but I have to use them every day with my small hands, switching out the generator battery.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: The bread machine seems to be slowly dying on us, and this is very sad.


Melanie Grande, Crew 186 Engineer

EVA Report – January 4th

Author: Cesare Guariniello

Purpose of EVA: Visit the regions of Beige Moon and Yellow Moon, collect samples of salts

Destination: Yellow Moon

UDM27 Coordinates:
515700 E
4254200 N

Participants: Cesare, Melanie, Mark, Samuel

Narrative: This EVA brought the crew for the first time Northwest of the habitat, in the region of The Moons (White Moon, Beige Moon, Gray Moon, Yellow Moon). After egressing the habitat airlock, the EVA crew members spent about fifteen minutes strapping and securing the spectrometer case and the geology toolbox to the ATVs. This simple activity was very instructive about the difficulties of performing simple tasks when donning bulky gloves and a space suit, being able to communicate only via radio, while always “thinking like a Martian” (for example, this means slow and accurate movements, paying attention to the safety of all crew members).

Due to the low temperature, the crew had to use the choker to start the ATVs. However, despite warming the 300, its engine quit multiple times. Since the EVA had multiple stops planned, the crew decided not to use the 300, and instead added a third 350 (vehicles used were the Honda, 350.1, 350.2, and 350.3). The ride North was quite slow, due to the presence of a delicate, expensive instrument.

Once again, the crew did not spot the unmarked road (in this case, Hwy 1572), and went further North to the end of the ATV road. This detour gave the crew the first spectacular sight: on the way back towards Hwy 1572, the crew drove through the Glistening Seas, and the position of the Sun showed how appropriate this name is. The whole plain was sparkly with reflections from gypsum crystals, which gave a sense of awe. In the middle of the Glistening Seas, the EVA crew spotted the location of Hwy 1572 and followed it towards Yellow Moon. Upon reaching the location and parking the ATVs, the crew geologist began his collection of samples, while other EVA crew members explored the area. The magnificent landscape, open to the East on the red Morrison formation, and shadowed to the West by the brown rims of Mancos shale, prompted a discussion within the EVA crew to discuss about the possibility of reaching the Martian Moons.

Due to the delays at the beginning of the EVA, the crew had to turn around after just 40 minutes in the field. On the way back, two more stops for collection of geological samples were added, before the crew stopped at Mount Nutella (see photos of the day) to climb the sand dune and look at the formations from the top of the mount. Melanie lost communication at this point, but this did not cause any problem: she was put in the middle of the ATVs convoy, and the EVA party was already enroute to the habitat. Upon returning, while Melanie refueled the ATVs and tested the 300 (which worked, since it had warmed up), and Cesare unstrapped the tools, the crew was welcomed by a new pet: the NorCal Mars Society’s rover Phoenix!! The EVA crew played with it for a little bit, before ingressing the airlock to end this astonishing EVA.

Journalist Report – January 4th

[Sol 4]

Today proved relatively restful for most of the team. Cesare Guariniello, our geologist, cooked chocolate pancakes for the crew after an early morning yoga session once again led by Kshitij Mall, the executive officer. Though a hearty EVA to the north in search of salt deposits was planned, it was cut short by unanticipated difficulties with one of the ATVs. The EVA team of 4 returned just after 1 pm and, given the strenuous expeditions of the last few days, was more than happy to spend the remainder of the afternoon catching up on various tasks. We deployed a small robotic rover built by the NorCal Mars Society and took turns operating it remotely in the vicinity of the hab. Mark Gee, the green hab scientist, has also made significant progress on his microgreens research and gave the team a tour of his experiments. The gesture complimented the lecture Cesare gave to us last night on the local geology.

Tonight the team will try our hand at making pizza out of the abundance of powdered and freeze dried ingredients. Afterwards we will continue to relax over card games and television shows. Hopefully it will be enough to ease tensions over the nutella. Today’s EVA crew ascended a ridge named “Mount Nutella” (no joke) but the effort proved futile in replenishing our reserves.

Justin Mansell, MDRS Crew 186 Journalist

Science Report – January 4th

Science: Geology

EVA #4 brought our crew to the Northwest of the habitat, in a region called “Yellow Moon”. The whole region sits between the top member of the Morrison formation (Brushy Basin) and the bottom of Mancos shale, and it is heavily weathered, with soft clay soil. The predominant geological feature of this region is the presence of salts, especially gypsum, from the region of Glistening Seas, through Beige Moon and Grey Moon, all the way to Yellow Moon. Besides samples of crystals, we collected some sandstones with traces of salts.
The different accretion of gypsum crystals, most often occurring in thin layers, and rarely in romboid shape, suggests different water and evaporation processes. The samples will be analyzed in the laboratory, since the short EVA time was just enough for in-situ selection and collection. The area presents outstanding views of the whole region, as described in the EVA report.

Thank you

GreenHab Report – January 4th

Mark Gee


Environmental control:

Cooling with fan, open door and vent


Shade cloth on crops

Working Hour: 06:00 PM
Inside temp at working hour: 17 C
Outside temp during working hours: -2 C
Inside temperature high: 34 C
Inside temperature low: 13 C
Inside humidity: 30 %RH
Inside humidity high: 31 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: Transplanted 20 tomato plants into large, individual pots. They will need cages soon. Cold air from the fan was blowing directly across the beans and causing them to shrivel. We built a baffle out of cardboard to redirect the cold air towards the ceiling, which creates better airflow and protects the plants. I have made space and plan to plant lettuce, carrots, kale, and sprouts.

Daily water usage for crops: 15 gallons

Time(s) of watering for crops: 09:30AM, 05:30 PM

Research observations: Microgreens sprouting in all trials. Moringa experiments still growing.

Changes to research plants: None.

Aquaponics: Not in use.

Narrative: Most of my time with the plants was spent transplanting tomatos becaus I’ve inherited over 50 tomato plants, previously planted four to a pot. It seems like someone got a tad excited when planting, but we have room for now.

My fellow crew members asked for a tour of the Green Hab, so I walked them through what we have growing, what is near harvest, how the environmental controls work, and what research projects are going on. Melanie especially liked the baby cucumbers which smaller than her fingernail. I told them about the challenges with environmental control and how the cold air from the fan was damaging the beans. Max came up with a clever idea for a baffle that directs the cold air upwards which creates better airflow and protects the plants.

The first two days after seeding an experiment I’m always worried that the plants won’t grow because the surface bleach seeped in and killed them, or there was too much water, or not enough water, or the temperature was wrong, or something else. I spend the next 48 hours checking in on my plants way too often and wondering if the lifeless brown spheres will unfold into leaves, roots, and data that can be published. Fortunately, the time of waiting is over. The seeds took their time, but now every single tray of microgreens planted has an abundance of sprouts that are growing with gusto.

Support/supplies needed: None